For this function, calculate the equation of the tangent line at the given point
f(x) = sqrt of 6x at the point (6,6)

Answer by tommyt3rd(5050)

df/dx=1/(2sqrt(6x))(6)=3/sqrt(6x)
at x=6, m=3/6=1/2

line:
y-6=1/2(x-6)